# Enabling Local Neural Operators to perform Equation-Free System-Level Analysis

### v1

@@ Results
Figure 5: RandONet for the parabolic Liouville–Bratu–Gelfand PDE in Eq. (21), trained using trajectories at a fixed value of the bifurcation parameter ($\lambda=1$). Results are compared with a Finite Difference (FD) based time-stepper. (a) Steady-state solutions obtained with Newton-GMRES; (b) Full spectrum of the steady-state Jacobian of the actual PDE time-stepper; in the blow up panel, the three leading eigenvalues as computed using the Arnoldi method.
@@
Finally, we trained a homotopy-based RandONet to learn the dynamics in the full range of the bifurcation parameter in $\lambda\in[0,3.8]$. For that purpose, we used $1500$ neurons in the branch and $150$ neurons in the trunk. In the trunk network, which processes low-dimensional spatial coordinates, we applied a sigmoidal transformation with randomized weights. These weights are sampled following the approach described in [31, 52, 48, 37]. Figure 6a shows the bifurcation diagram obtained wrapping around the RandONet, Newton-GMRES with pseudo-arclength continuation. As shown, the bifurcation diagram obtained with a RandONet is, for any practical means, identical up to the tolerance of $10^{-5}$, to the reference one, accurately capturing both stable and unstable branches. The saddle-node bifurcation was identified at $\lambda_{c}=3.51381$, a value that corresponds to the theoretical one with an error of the order $10^{-5}$. Figure 6b shows the approximation of the real part of the three leading eigenvalues as a function of $\lambda$ as obtained by RandONets by applying the Arnoldi method. As shown, RandONets approximate the actual eigenvalues and stability of the solution with high accuracy.
Figure 6: RandONet for the Liouville-Bratu-Gelfand PDE in Eq. (21), trained using the full dataset of the dynamics in $\lambda\in[0,3.8]$. (a) Reconstructed bifurcation diagram with the RandONets, compared with the true FD bifurcation diagram. In the two insets, we show two representative solutions associated with $\lambda=1$ (stable branch) and $\lambda=2$ (unstable branch). (b) Tracking the absolute value of first 3 leading eigenvalues.
@@
Here, we illustrate how we can further accelerate time simulations by applying projective integration (PI) using the RandONet time-stepper with a small time step ($dt=0.001$). Projective integration (for problems with time-scale separation) alternates between short relaxation steps and large projective steps to advance efficiently along the slow time-evolution manifold. Specifically, we performed five short relaxation steps with $\Delta t=0.001$ (total relaxation time $0.005$), followed by a large projective step of size $\Delta T=0.015$ using a backward Euler data-driven estimation of the solution time derivative. Figure 7 summarizes the performance of the RandONet forward Euler projective integrator. Figure 7a shows the evolution of the PI solution $u(x,t)$, and Figure 7b displays the related space-time error. This error remains confined within $10^{-3}$, demonstrating excellent agreement with the reference solver. Figures 7c-7e track the evolution of the $L^{\infty}$-error in the solution and its first and second spatial derivatives over time. The first derivative error, representing total variation, and the second derivative error, representing total curvature, highlight the relaxation process. Notably, as shown in Figure 7e there is a rapid decay in total curvature, confirming the fast convergence to the slow manifold. This behavior reflects the typical separation of timescales in dissipative PDEs, where higher spatial derivatives correspond to higher frequency modes, which tend to relax more rapidly than the field component $u$.
@@
Finally, we demonstrate how a NO can be used to learn and exploit a local in space solution operator within the “Gap-Tooth“ framework. In particular, we discretized spatial domain of the Liouville–Bratu–Gelfand PDE into $N_{\text{teeth}}=21$ teeth and $N_{\text{gaps}}=20$ gaps of equal width, so that the input functions are defined “inside the tooth,” evolving over smaller spatial domains. Thus, they require smoother, less variegated families of initial conditions. This local‐in‐space setup avoids full‐domain coverage (here, coupled local NOs only evolve over half of the domain, and the other half –the gaps– are recovered by smoothness/interpolation); this enables more data-efficient training, and accelerates EF computations. For approximating the Gap-Tooth solution operator, we used a RandONet with $300$ neurons in the branch and $100$ in the trunk. We obtain results in a mean squared error (MSE) of $1.37E$-$11$, a median L2 error of $5.13E$-$06$, and a maximum absolute error of $9.18E$-$05$. Notably, the training of the single-parameter “Gap-Tooth” RandONet took $0.15$ seconds. We provide further details on the data generation and training process in Appendix sections C.3 and C.4. We used the forward Euler method to integrate the “Gap-Tooth” RandONet. Figure 8 compares the time evolution between the “Gap-Tooth” RandONet time stepper and the reference “Gap-Tooth” solver. Figure 8a shows the predicted evolution. As shown in Figure 8b, this evolution matches well with the reference trajectories. Figure 8c compares both steady-state solutions, again demonstrating that the “Gap-Tooth” RandONet captures the steady-state behavior accurately.
Spectral analysis of the “Gap-Tooth” scheme shown in Figures 8d reveals that the three leading eigenvalues match very closely with the full reference problem solution ones.
(a) Gap-Tooth NO timestepper
(b) error
(c) steady-state
(d) full spectrum                            leading eigenvalues (Arnoldi)
Figure 8: Gap-Tooth RandONet timestepper results for the Liouville-Bratu-Gelfand PDE in Eq. (21), with $\lambda=1$. (a) Comparison of the Gap-Tooth RandONet with the reference finite differences with forward Euler time stepper; (b) Corresponding temporal-spatial error; (c) Steady states of both time steppers computed with Jacobian-Free Newton-Krylov; (d) Full eigenvalue spectrum of the RandONet and reference Jacobians in their respective steady states; in the zoom in, the 3 leading eigenvalues computed with the Arnoldi method.
@@
To generate the training data for RandONet, we employed the $D1Q3$-Lattice-Boltzmann method (LBM) [7, 5, 71, 72] to simulate the spatiotemporal dynamics through particles with discrete velocities, as detailed in Appendix section D.1. We used random sigmoid-like initial conditions, and we considered bifurcation parameter values in the range $\varepsilon\in[-0.005,0.955]$. The lower value represents stable oscillating solutions beyond the Hopf bifurcation, while the upper value is the location of the saddle-node bifurcation.
@@ Discussion
In this work, we have demonstrated the application of local NOs – local in time, over short-time windows, and possibly local in space, over small spatial patches – to perform system-level analysis, with a particular focus on long-time steady-state, stability, and bifurcation analysis. Through the use of advanced methods such as homotopy-based embedding NOs (for our illustrations RandONets) and the exploration of multiscale techniques like Projective Integration (PI) and “Gap-Tooth” schemes, we have shown how scientific machine learning can be integrated with traditional numerical analysis methods (in the absence of an explicit, closed-form equation) to offer efficient, scalable solutions to complex, spatiotemporally varying dynamical systems.
@@
In this work, we performed some illustrative comparison between “vanilla” DeepONets and their similar but shallow and randomized counterparts – “vanilla” RandONets – specifically on the Allen-Cahn equation. For the Liouville–Bratu–Gelfand PDE and the FitzHugh–Nagumo equations, only RandONets were used, as they exhibited favorable performance and training efficiency in the first benchmark. A detailed comparison with alternative NO architectures – such as DeepONets and Fourier neural operators – for learning PDE solution operators (including higher-dimensional PDEs and hyperbolic systems with conservation laws) remains an important direction for future research. More broadly, our key message is that every variant of neural operator benefits from the combination of local training and classical numerical analysis workflows, regardless of the specific implementation or example.
@@
It has already been argued (in [54]) that iterating short time neural operators (trained in their case not on data, but in an entirely self-supervised manner as physics informed DeepONets) exhibits advantages over long-time PINN training. In our case, we argue that the use of such local in time (and possibly in space) NOs, trained either on data or in a physics-informed manner, opens up exciting algorithmic possibilities. We anticipate that an important benefit – in the case of data-driven NOs – is the potential to “get away with” much more small and parsimonious training data sets. Indeed, we expect that it is sufficient to train “patch” NOs with relatively low-dimensional families of initial and boundary conditions (constant, linear, possibly quadratic) - much less variegated than large-space initial conditions and long-time boundary conditions (and, for that matter, different domain shapes). We also expect, in the spirit of [54]) that entirely self-supervised physics-informed NOs will also be easier to train if they are local-in-space $\times$ time.
@@
Finally, we have shown that our neural operators can also be used “inside” Projective Integration (PI) (as well as telescopic projective integration, TPI [59]) schemes. PI can significantly accelerate transient simulations as long as the “inner” time-stepper can correctly dampen the fast transient modes. We demonstrated that a Local Neural Operator can be seamlessly included in such a framework.
In current (and future) work, we extend our implementation to include adaptive time stepping, and to combine Projective Integration with the “Gap-Tooth” scheme to obtain solutions accelerated by doubly local (in space and in time) NOs. It is important to note that, with local NOs playing the role of discretization points in traditional numerical schemes – with the “right” flux communication between them, as prescribed by numerical analysis – our approach can also be “geometry-agnostic”. Indeed, to the extent that different macroscale domains can be usefully modeled through point – or finite volume –discretizations, local patch NOs hold the promise of liberating a class of NO computations from needing to be trained on data from specific geometric domains.
@@ Methods
Data and code used to produce the findings of this study are publicly available on the GitHub repository https://github.com/Centrum-IntelliPhysics/Time-stepper.
@@ Acknowledgments
G.F., H.V., S.G., and I.G.K. acknowledge the Department of Energy (DOE) support under Grant No. DE-SC0024162, as well as partial support by the National Science Foundation. The authors would like to acknowledge computing support provided by the Advanced Research Computing at Hopkins (ARCH) core facility at Johns Hopkins University and the Rockfish cluster. ARCH core facility (rockfish.jhu.edu) is supported by the National Science Foundation (NSF) grant number OAC1920103. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the funding organizations. C.S. acknowledges partial support by the PNRR MUR projects PE0000013-Future Artificial Intelligence Research-FAIR & CN0000013 CN HPC - National Centre for HPC, Big Data and Quantum Computing, Gruppo Nazionale Calcolo Scientifico-Istituto Nazionale di Alta Matematica (GNCS-INdAM).
@@ B.1 Data generation for the Allen-Cahn Equation
The Allen-Cahn PDE in Eq. (14) is solved numerically to generate training and test data for different values of the bifurcation parameter $\varepsilon$. We simulate the PDE using the chebfun [67, 68] package in Matlab, which uses a pseudospectral Chebyshev method for spatial discretization of the PDE [67, 68], combined with a method of lines approach. We further use implicit time stepping of variable order and variable step size as implemented in the chebfun suite pde15s solver, an expansion of the well-known ode15s [69]. The boundary conditions are set to Neumann-type, imposing $\partial u/\partial x=0$ at $x=\pm 1$. We vary the parameter $\varepsilon$ between $0.22$ and $0.7$ across $N_{\text{par}}=25$ evenly spaced values.
@@
  • Flat and Zero Solutions: In addition to the sine and cosine modes, initial conditions also include zero solutions, which represent the unstable steady state, and flat solutions with random perturbations in the range $[-1.5,1.5]$ to explore the solution space more broadly.
@@
For each $\varepsilon$ and initial condition, the solution $u(x,t)$ is computed and stored at all time steps. We duplicate the dataset with symmetric solutions $-u(x,t)$ due to the symmetric structure of the Allen-Cahn equation. The dataset is then organized into training and test sets, with:
@@ C.2 Training of the RandONet Neural Operators for the Liouville-Bratu-Gelfand PDE
In Matlab, we train RandONet architectures on the training data for the Liouville-Bratu-Gelfand PDE and we evaluate its performance. The process is set up for both the parametric and non-parametric scenarios. In particular, the number of neurons in the trunk layer is set to $150$. We train RandONets for a different number of neurons in the branch hidden layer: $[10,20,40,80,100,150,300,500,750,1000,1500,2000]$. Our code investigates the convergence of the method with respect to these neuron counts in the branch net.

### v2

@@ Results
Figure 6: RandONet for the parabolic Liouville–Bratu–Gelfand PDE in Eq. (21), trained using trajectories at a fixed value of the bifurcation parameter ($\lambda=1$) (in (a)-(b)) and trained using the full dataset of the dynamics in $\lambda\in[0,3.8]$ (in (c)-(d)). Results are compared with a Finite Difference (FD) based time-stepper. (a) Steady-state solutions obtained with Newton-GMRES; (b) Full spectrum of the steady-state Jacobian of the actual PDE time-stepper; in the blow up panel, the three leading eigenvalues as computed using the Arnoldi method. (c) Reconstructed bifurcation diagram with the RandONets, compared with the true FD bifurcation diagram. In the two insets, we show two representative solutions associated with $\lambda=1$ (stable branch) and $\lambda=2$ (unstable branch). (d) Tracking the absolute value of first 3 leading eigenvalues.
@@
Finally, we trained a homotopy-based RandONet to learn the dynamics in the full range of the bifurcation parameter in $\lambda\in[0,3.8]$. For that purpose, we used $1500$ neurons in the branch and $150$ neurons in the trunk. In the trunk network, which processes low-dimensional spatial coordinates, we applied a sigmoidal transformation with randomized weights. These weights are sampled following the approach described in [31, 66, 48, 37]. Figure 6c shows the bifurcation diagram obtained wrapping around the RandONet, Newton-GMRES with pseudo-arclength continuation. As shown, the bifurcation diagram obtained with a RandONet is, for any practical means, identical up to the tolerance of $10^{-5}$, to the reference one, accurately capturing both stable and unstable branches. The saddle-node bifurcation was identified at $\lambda_{c}=3.51381$, a value that corresponds to the theoretical one with an error of the order $10^{-5}$. Figure 6d shows the approximation of the real part of the three leading eigenvalues as a function of $\lambda$ as obtained by RandONets by applying the Arnoldi method. As shown, RandONets approximate the actual eigenvalues and stability of the solution with high accuracy.
@@
Here, we illustrate how we can further accelerate time simulations by applying projective integration (PI) using the RandONet time-stepper with a small time step ($dt=0.001$). Projective integration (for problems with time-scale separation) alternates between short relaxation steps and large projective steps to advance efficiently along the slow time-evolution manifold.
@@
Specifically, we performed five short relaxation steps with $\Delta t=0.001$ (total relaxation time $0.005$), followed by a large projective step of size $\Delta T=0.015$ using a backward finite difference estimation of the solution time derivative. Figure 7 summarizes the performance of the RandONet forward Euler projective integrator. Figure 7a shows the evolution of the PI solution $u(x,t)$, and Figure 7b displays the related space-time error. This error remains confined within $10^{-3}$, demonstrating excellent agreement with the reference solver. Figures 7c-7e track the evolution of the $L^{\infty}$-error in the solution and its first and second spatial derivatives over time. The first derivative error, representing total variation, and the second derivative error, representing total curvature, highlight the relaxation process. Notably, as shown in Figure 7e there is a rapid decay in total curvature, confirming the fast convergence to the slow manifold. This behavior reflects the typical separation of timescales in dissipative PDEs, where higher spatial derivatives correspond to higher frequency modes, which tend to relax more rapidly than the field component $u$.
(a) Gap-Tooth NO timestepper
(b) Gap-Tooth NO error
(c) steady-state
(d) Patch NO timestepper
(e) Patch NO error
(f) full spectrum           leading eigenvalues (Arnoldi)
Figure 8: Gap-Tooth RandONet timestepper results for the Liouville-Bratu-Gelfand PDE in Eq. (21), with $\lambda=1$. (a) the Gap-Tooth RandONet timestepper; (b) Corresponding spatio-temporal error w.r.t the reference solution; (c) Patch RandONet timestepper (Combination of Gap-Tooth NO and PI); (d) correspponding spatio-temporal error w.r.t the reference solution; (e) Steady states of both time steppers computed with Jacobian-Free Newton-Krylov; (f) Full eigenvalue spectrum of the RandONet and reference Jacobians in their respective steady states; in the zoom in, the 3 leading eigenvalues computed with the Arnoldi method. (e)-(f) results are compared with a Finite Difference (FD) based time-stepper.
@@
Finally, we demonstrate how a NO can be used to learn and exploit a local in space solution operator within the “Gap-Tooth“ framework. In particular, we discretized spatial domain of the Liouville–Bratu–Gelfand PDE into $N_{\text{teeth}}=21$ teeth and $N_{\text{gaps}}=20$ gaps of equal width, so that the input functions are defined “inside the tooth,” evolving over smaller spatial domains. Thus, they require smoother, less variegated families of initial conditions. This local‐in‐space setup avoids full‐domain coverage (here, coupled local NOs only evolve over half of the domain, and the other half –the gaps– are recovered by smoothness/interpolation); this enables more data-efficient training, and accelerates EF computations. For training and short-time predictions, we used sampled data with a local time step of $\Delta t=0.0001$. For approximating the Gap-Tooth solution operator, we used a RandONet with $300$ neurons in the branch and $100$ in the trunk. We obtain results in a mean squared error (MSE) of 1.37E1.37E-$11$, a median L2 error of 5.13E5.13E-$06$, and a maximum absolute error of 9.18E9.18E-$05$. Notably, the training of the single-parameter “Gap-Tooth” RandONet took $0.15$ seconds. We provide further details on the data generation and training process in Appendix sections C.3 and C.4. We used the forward Euler method to integrate the “Gap-Tooth” RandONet. Figure 8 compares the time evolution between the “Gap-Tooth” RandONet time stepper and the reference “Gap-Tooth” solver. Figure 8a shows the predicted evolution. As shown in Figure 8b, this evolution matches well with the reference trajectories. Figure 8c compares both steady-state solutions, again demonstrating that the “Gap-Tooth” RandONet captures the steady-state behavior accurately.
Spectral analysis of the “Gap-Tooth” scheme shown in Figures 8f reveals that the three leading eigenvalues match very closely with the full reference problem solution ones.
The “patch” NO
We further extend the Gap-Tooth NO approach by incorporating Projective Integration, resulting in a Patch Neural Operator (Patch NO). This combines local-in-space and local-in-time evolution, enhancing both computational efficiency. Specifically, we performed fifty short relaxation steps with $\Delta t=0.0001$ (total relaxation time $0.005$), followed by a large projective step of size $\Delta T=0.015$ using a backward finite-difference estimation of the solution time derivative. Figures 8d and 8e show the Patch NO timestepper and the corresponding error. This error remains confined within $10^{-3}$, demonstrating excellent agreement with the reference solver. As in the Gap-Tooth case, the Patch NO accurately reproduces the reference dynamics while reducing temporal resolution requirements through projective updates.
@@
To generate the training data for RandONet, we employed the $D1Q3$-Lattice-Boltzmann method (LBM) [7, 5, 82, 83] to simulate the spatiotemporal dynamics through particles with discrete velocities, as detailed in Appendix section D.1. In particular, a local time step of $\Delta t=0.01$ was used during the simulation, and data snapshots were collected every $\Delta t=0.1$ for training and short-time predictions of the NO. For longer time predictions, the NO was applied autoregressively. We used random sigmoid-like initial conditions, and we considered bifurcation parameter values in the range $\varepsilon\in[0.005,0.955]$. The lower value represents stable oscillating solutions beyond the Hopf bifurcation, while the upper value is the location of the saddle-node bifurcation.
@@ Discussion
In this work, we have demonstrated the application of local NOs – local in time, over short-time windows, and possibly even local in space, over small spatial patches – to perform system-level analysis, with a particular focus on long-time steady-state, stability, and bifurcation analysis. Through the use of advanced methods such as homotopy-based embedding NOs (for our illustrations RandONets) and the exploration of multiscale techniques like Projective Integration (PI) and “Gap-Tooth” schemes, we have shown how scientific machine learning can be integrated with traditional numerical analysis methods (in the absence of an explicit, closed-form equation) to offer efficient, scalable solutions to complex, spatiotemporally varying dynamical systems.
@@
In this work, we performed some illustrative comparison between “vanilla” DeepONets and their similar but shallow and randomized counterparts – “vanilla” RandONets – specifically on the Allen-Cahn equation. For the Liouville–Bratu–Gelfand PDE and the FitzHugh–Nagumo equations, only RandONets were used, as they exhibited favorable performance and training efficiency in the first benchmark. A detailed comparison with alternative NO architectures – such as DeepONets and Fourier neural operators – for learning PDE solution operators (including higher-dimensional PDEs and hyperbolic systems with conservation laws) remains an important direction for future research. While such architectures have demonstrated promise, they also present notable challenges. DeepONets and FNOs, in particular, often involve millions of parameters and require costly, high-dimensional non-convex optimization, which can hinder both training efficiency and generalization. In our experiments, these difficulties manifested, for example, in the Allen–Cahn benchmark, where the DeepONet-based timestepper struggled to yield reliable results for stability analysis. To mitigate such issues, potential improvements can include Sobolev training to better capture higher-frequency information induced by derivatives [54, 84, 85], and the enforcement of dissipation to steer out-of-distribution predictions toward attractors [54]. These directions remain exploratory, but they highlight the need for physically informed constraints in complex operator learning. More broadly, our key message is that every variant of neural operator – including DeepONet and FNO – can benefit from the combination of local training and classical numerical analysis workflows, regardless of the specific implementation or example.
@@
It has already been argued (in [53]) that iterating short time neural operators (trained in their case not on data, but in an entirely self-supervised manner as physics informed DeepONets) exhibits advantages over long-time PINN training. In our case, we argue that the use of such local in time (and possibly in space) NOs, trained either on data or in a physics-informed manner, opens up exciting algorithmic possibilities. We anticipate that an important benefit - in the case of data-driven NOs - is the potential to “get away with” much more small and parsimonious training data sets. Indeed, we expect that it is sufficient to train “patch” NOs with relatively low-dimensional families of initial and boundary conditions (constant, linear, possibly quadratic) - much less variegated than large-space initial conditions and long-time boundary conditions (and, for that matter, different domain shapes). We also expect, in the spirit of [53]), that entirely self-supervised physics-informed NOs will also be easier to train if they are local-in-space $\times$ time.
@@
Finally, we have shown that our neural operators can also be used “inside” Projective Integration (PI) (as well as telescopic projective integration [56]) schemes. PI can significantly accelerate transient simulations as long as the “inner” time-stepper can correctly dampen the fast transient modes. We demonstrated that a Local Neural Operator can be seamlessly included in such a framework.
In current (and future) work, we extend our implementation to include adaptive time stepping, and to combine Projective Integration with the “Gap-Tooth” scheme to obtain solutions accelerated by doubly local (in space and in time) NOs. It is important to note that, with local NOs playing the role of discretization points in traditional numerical schemes – with the “right” flux communication between them, as prescribed by numerical analysis – our approach can also be “geometry-agnostic”. Indeed, to the extent that different macroscale domains can be usefully modeled through point, or finite volume, discretizations, local patch NOs hold the promise of liberating a class of NO computations from needing to be trained on data from specific geometric domains.
@@ Methods
Data and code used to produce the findings of this study are publicly available on the GitHub repository local-neural-operator-time-stepper.
@@ Acknowledgments
G.F., H.V., S.G., and I.G.K. acknowledge the Department of Energy (DOE) support under Grant No. DE-SC0024162; I.G.K.also acknowledges partial support by the National Science Foundation under Grants No. CPS2223987 and FDT2436738. The authors would like to acknowledge computing support provided by the Advanced Research Computing at Hopkins (ARCH) core facility at Johns Hopkins University and the Rockfish cluster. ARCH core facility (rockfish.jhu.edu) is supported by the National Science Foundation (NSF) grant number OAC1920103. C.S. acknowledges partial support by the PNRR MUR projects PE0000013-Future Artificial Intelligence Research-FAIR & CN0000013 CN HPC - National Centre for HPC, Big Data and Quantum Computing, Gruppo Nazionale Calcolo Scientifico-Istituto Nazionale di Alta Matematica (GNCS-INdAM). Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the funding organizations.
@@ B.1 Data generation for the Allen-Cahn Equation
The Allen-Cahn PDE in Eq. (14) is solved numerically to generate training and test data for different values of the bifurcation parameter $\varepsilon$. We simulate the PDE using the chebfun [78, 79] package in Matlab, which uses a pseudo-spectral Chebyshev method for spatial discretization of the PDE [78, 79], combined with a method of lines approach. We further use implicit time stepping of variable order and variable step size as implemented in the chebfun suite pde15s solver, an expansion of the well-known ode15s [80]. The boundary conditions are set to Neumann-type, imposing $\partial u/\partial x=0$ at $x=\pm 1$. We vary the parameter $\varepsilon$ between $0.22$ and $0.7$ across $N_{\text{par}}=25$ evenly spaced values.
@@
  • Flat and Zero Solutions: In addition to the sine and cosine modes, initial conditions also include zero solutions, which represent the unstable steady state, and flat solutions with random perturbations in the range [1.5,1.5][-1.5,1.5] to explore the solution space more broadly.
@@
For each $\varepsilon$ and initial condition, the solution $u(x,t)$ is computed and stored at all time steps. We duplicate the dataset with symmetric solutions u(x,t)-u(x,t) due to the symmetric structure of the Allen-Cahn equation. The dataset is then organized into training and test sets, with:
@@ C.2 Training of the RandONet Neural Operators for the Liouville-Bratu-Gelfand PDE
In Matlab, we train RandONet architectures on the training data for the Liouville-Bratu-Gelfand PDE and we evaluate its performance. The process is set up for both the parametric and non-parametric scenarios. In particular, the number of neurons in the trunk layer is set to $150$. We train RandONets for a different number of neurons in the branch hidden layer: [10,20,40,80,100,150,300,500,750,1000,1500,2000][10,20,40,80,100,150,300,500,750,1000,1500,2000]. Our code investigates the convergence of the method with respect to these neuron counts in the branch net.